# No Minima, No Collisions: Combining Modulation and Control Barrier Function Strategies for Feasible Dynamical Collision Avoidance

### v1

@@ VI-A Speed-Constraining Mod-DS
When the unconstrained output from Mod-DS $u_{\text{unc}}=\dot{x}_{\text{unc}}$ exceeds the speed limit, solving the following Quadratically Constrained Linear Programming (QCLP) problem finds the constrained output $u_{\text{c}}$ that is most similar to $u_{\text{unc}}$ and ensures safety (42) while satisfying speed input constraint $u_{\text{ub}}$ (41). In both speed-constraining and velocity-constraining Mod-DS, robot safety is ensured by restricting the robot’s velocity moving towards the obstacle, i.e. the projection of $u_{\text{c}}$ onto vector $n(x,t)$, to be no larger than that of $u_{\text{unc}}$.
@@
$\langle n,u\rangle\geq\min(\langle n,\dot{x}_{o}\rangle,\langle n,u_{\text{unc}}\rangle)$ (42)
@@
$n^{\top}u\geq\min(n^{\top}\dot{x}_{o},n^{\top}u_{\text{unc}})$
Denote $v_{n}=\min(n^{\top}\dot{x}_{o},n^{\top}u_{\text{unc}})$. Solving for the explicit solutions of the above convex optimization problem using Karush–Kuhn–Tucker (KKT) conditions, a simple explicit solution is found as in (44), where $v_{e}=\sqrt{||u_{\text{unc}}||_{2}^{2}-||n^{\top}u_{\text{unc}}||_{2}^{2}}=||H^{\top}u_{\text{unc}}||_{2}$ is the projection of $u_{\text{unc}}$ onto the hyperplane $H=H(x,t)$ tangent to the boundary function $h(x,t)$ (16).
@@ VI-C Theoretical Guarantees
Consider an obstacle whose boundary is defined as $h(x,t)=0$. Any trajectory $x(t)$, $t\in[0,\infty)$ starting outside the obstacle boundary, i.e., $h(x(0),0)\geq 0$, evolving according to Eq. (14) and being constrained by the Speed-Constraining and Velocity-Constraining approaches Mod-DS presented in Section VI-A and VI-B, will never penetrate the obstacle boundary, i.e., $h(x(t),t)\geq 0$, $\forall t\in[0,\infty)$.
@@
VII Modulation-Based Control Barrier Functions (MCBF)
In fully actuated systems, Mod-DS approaches outperform CBF-QP in navigating around non-convex unsafe sets with far fewer undesirable local minima. However, CBF-QP is still more commonly used in robot applications because it works for control affine systems in general and the QP formulations are more intuitive to the users. Nevertheless, choosing between superiority in handling complex robot dynamics and navigating among concave unsafe sets free of undesirable local minima should not be a tradeoff that users have to make when designing robot controllers. In this section, 2 Mod-based CBF-QP (MCBF-QP) approaches that combine the strength of CBF-QP with Mod-DS are proposed to ensure respectively few-local-minimum and local-minimum-free navigation in concave-obstacle environments with the ease of incorporating different robot dynamics.
@@ VII-A Reference MCBF for Star-shaped Obstacle Avoidance
In subsection V-C, theoretical analysis of reference Mod-DS and CBF-QP demonstrated that CBF-QP’s performance in $n(x,t)$ direction is similar to that of the reference Mod-DS in Theorem V.5. Besides, in Theorem V.6, reference Mod-DS’s introduction of off-diagonal components in the tangent hyperplane $H(x,t)$ navigation is identified to be the major differences between CBF-QP and reference Mod-DS in fully-actuated systems and the key factor for reducing undesirable equilibrium in star-shaped obstacle avoidance. Adding additional constraints for $u_{\text{cbf}}$ on the tangent hyperplane will reduce undesirable equilibrium caused by CBF constraints. In theory, there are numerous ways to incorporate the off-diagonal components into CBF-QP results. Here, we propose a linear constraint formulation that is mathematically simple and does not increase the complexity of the original CBF-QP problem. Note that $E_{r}(x,t)^{-1}[2:d,:]\dot{x}=H(x,t)^{\top}(I-\frac{1}{\nabla_{x}h(x,t)^{\top}r}r\nabla_{x}h(x,t)^{\top})\dot{x}$ solves for the projection of $\dot{x}$ onto hyperplane $H(x,t)$ in basis $E_{r}(x,t)$. $g(x)$ and $E_{r}(x,t)^{-1}$ are defined in (8) and (32).
@@
$L_{f}h(x,t)+L_{g}h(x,t)u+\frac{\partial h(x,t)}{\partial t}\geq-\alpha(h(x,t))$ (46)
$H(x,t)^{\top}(I-\frac{r(x,t)\nabla_{x}h(x,t)^{\top}}{\nabla_{x}h(x,t)^{\top}r(x,t)})g(x)(u-u_{\text{nom}})=\rho$ (47)
@@
$\nabla_{x}h(x,t)^{\top}u+\frac{\partial h(x,t)}{\partial t}\geq-\alpha(h(x,t))$ (49)
$H(x,t)^{\top}(I-\frac{r(x,t)\nabla_{x}h(x,t)^{\top}}{\nabla_{x}h(x,t)^{\top}r(x,t)})(u-u_{\text{nom}})=\rho$ (50)
Reference-MCBF-QP, compared with regular CBF-QP, improves the performance of the optimization-based safe controller for star-shaped unsafe sets by reducing local minimum numbers. The new method ensures that the safe control outputs will never lead to undesirable equilibria, except when $\dot{x}_{\text{nom}}$ is inversely collinear with the reference direction $r(x,t)$. The added constraints are designed to be of linear equality types and, thus would not cause large runtime increases. Additionally, the employment of relaxation parameter $\rho$ sustains the feasibility of the regular CBF-QP, i.e. given the same $x,t$, Reference-MCBF-QP will be feasible if and only if the CBF-QP without constraints (47) and (50) is feasible. The performance of fully-actuated reference MCBF-QP in star-shaped obstacle environments is validated in Figure 7.
@@
Derivation: Here we abuse notation and let $\alpha=\alpha(h(x,t))-\frac{\partial h(x,t)}{\partial t}$, $H=H(x,t)$, $r=r(x,t)$, and $\nabla_{x}h=\nabla_{x}h(x,t)$. MCBF-QP in (48) can be reformulated into Lagrangian function in (51), where $\mu_{\text{eq}}\in\mathbb{R}^{d-1}$ and $\mu_{\text{cbf}}$ are the Lagrange multipliers.
@@
$V=\begin{cases}\begin{bmatrix}u_{\text{nom}}\\ 0\end{bmatrix}\quad\text{if}\quad\nabla_{x}h^{\top}u_{\text{nom}}+\alpha\geq 0\\ \begin{bmatrix}F^{-1}(I+\nabla_{x}hA)&-A^{\top}\\ A&-\frac{AF\nabla_{x}h}{\nabla_{x}h^{\top}\nabla_{x}h}\end{bmatrix}\begin{bmatrix}Fu_{\text{nom}}\\ -\alpha\end{bmatrix}\quad\text{otherwise}\end{cases}.$ (55)
In other words, the explicit solution $u_{\text{mcbf}}$ becomes
$u_{\text{mcbf}}=\begin{cases}u_{\text{nom}}\quad\text{if}\quad\nabla_{x}h^{\top}u_{\text{nom}}+\alpha\geq 0\\ F^{-1}(I+\nabla_{x}hA)Fu_{\text{nom}}+A^{\top}\alpha\quad\text{otherwise}\end{cases},$ (56)
@@
From 5(b), it is noticeable that vector $r(x,t)$ can be represented as a weighted vector sum of $n(x,t)$ and a vector $\sum_{i=1}^{d-1}w_{i}e_{i}$ parallel to the hyperplane $H(x,t)$ (58), $w_{i}\in\mathbb{R}$. Like before, we denote $e_{i}=e_{i}(x,t)$ and $n=n(x,t)$ to simplify the math. Note that $\sum_{i=0}^{d-1}(w_{i})^{2}=1$ because by definition, $||r||_{2}=1$.
@@
Based on the relationship between $n$ and $\nabla_{x}h$ given in (17), $\nabla_{x}h=||\nabla_{x}h||_{2}n$. Plug the new representation of $r$ and $\nabla_{x}h$ into (54), $F$ can be orthogonally diagonalized as the following, where $E=E(x,t)$ is defined in (16) with $d(x,t)$ in (17).
@@
Remark: The explicit solution of regular CBF-QP in fully-actuated systems (13) can be rewritten using $n$ as the following. Note that $nn^{\top}u_{\text{nom}}=n^{\top}u_{\text{nom}}n$ for all $n,u_{\text{nom}}\in\mathbb{R}^{d}$.
@@
Since $\lambda$, $\lambda_{e}$ can be any real number satisfying (18), there exist an infinite amount of $\lambda$, $\lambda_{e}$ meeting the requirement $\lambda_{e}-\lambda=\frac{1}{2}$. Therefore, the conclusion can be drawn that there always exist feasible $\lambda$, $\lambda_{e}$ pairs such that outputs from reference MCBF-QP equal that from reference Mod-DS on boundaries of the unsafe sets. This guarantees that reference MCBF-QP is free of undesirable equilibria except when $u_{\text{nom}}$ is inversely collinear to $r(x,t)$ in the boundary set $\partial C$.
@@ VII-B On-Manifold MCBF for Non-star-shaped Obstacle Mod- Avoidance
By keeping CBF constraints in regular CBF-QP to account for Mod-DS’s policies in $n(x,t)$ direction and introducing new constraints (47) and (50) to mimic reference Mod-DS’s actions in the tangent hyperplane $H(x,t)$, reference MCBF-QP successfully combines the strength of both and realizes star-shaped obstacle avoidance using QP-based methods. While on-manifold Modulation, functioning under the aggregate effects of 3 policies, is too structurally complicated to be explicitly connected to CBF-QP on the theory level, its main component, like other Mod-DS approaches, modulates $\dot{x}$ in $n(x,t)$ direction similar to CBF constraints and prevents the formation of undesirable equilibrium by ordering the directions of $\dot{x}$ projected onto the tangent hyperplane $H(x,t)$ using $\phi(x,t)$. Therefore, by introducing $H(x,t)$ related constraints, inspired by $\phi(x,t)$ in on-manifold Mod-DS, into CBF-QP, on-manifold MCBF-QP is constructed to achieve local-minimum-free safe control for control affine systems in general (VII-B). The parameter $\gamma$ is a user-defined positive real number, i.e. $\gamma\in\mathbb{R}^{+}$. Unlike reference MCBF-QP, which stays feasible as long as the CBF-QP without constraint (47) is feasible, on-manifold MCBF-QP’s feasible domain is always smaller than that of the CBF-QP without constraint (71). Additionally, the larger $\gamma$ is, the smaller the feasible domain of on-manifold MCBF-QP will be. The performance of on-manifold MCBF-QP in concave obstacle environments is validated in Figure 8.
@@
$L_{f}h(x,t)+L_{g}h(x,t)u+\frac{\partial h(x,t)}{\partial t}\geq-\alpha(h(x,t))$ (70)
$\phi(x,t)^{\top}f(x)+\phi(x,t)^{\top}g(x)u\geq\gamma$ (71)
@@
$\nabla_{x}h(x,t)^{\top}u+\frac{\partial h(x,t)}{\partial t}\geq-\alpha(h(x,t))$ (72)
$\phi(x,t)^{\top}u\geq\gamma$ (73)
@@
Remark: While reference MCBF-QP can be in theory extended to nonlinear dynamical systems with equations in (VII-A), its applications in high dimensional systems is limited because finding reference directions $r(x,t)$ when $d>3$ is nontrivial in general and its complexity scales up as $d$ increases. However, given proper step size $\beta$, the geodesic approximation method in (22) is able to closely approximate unsafe set boundaries for any given state $x\in\mathbb{R}^{d}$. As a result, on-manifold MCBF-QP can be readily applied in any control affine robot system.

### v2

@@ VI-A Speed-Constraining Mod-DS
When the unconstrained output from Mod-DS $u_{\text{unc}}=\dot{x}_{\text{unc}}$ exceeds the speed limit, solving the following Quadratically Constrained Linear Programming (QCLP) problem finds the constrained output $u_{\text{c}}$ that is most similar to $u_{\text{unc}}$ and ensures safety (46) while satisfying speed input constraint $u_{\text{ub}}$ (45). In both speed-constraining and velocity-constraining Mod-DS, robot safety is ensured by restricting the robot’s velocity moving towards the obstacle, i.e. the projection of $u_{\text{c}}$ onto vector $n(x,x_{o})$, to be no larger than that of $u_{\text{unc}}$.
@@
$\langle n,u\rangle\geq\min(\langle n,\bar{\dot{x}}_{o}\rangle,\langle n,u_{\text{unc}}\rangle)$ (46)
@@
$n^{\top}u\geq\min(n^{\top}\bar{\dot{x}}_{o},n^{\top}u_{\text{unc}})$
Denote $v_{n}=\min(n^{\top}\bar{\dot{x}}_{o},n^{\top}u_{\text{unc}})$. Solving for the explicit solutions of the above convex optimization problem using Karush–Kuhn–Tucker (KKT) conditions, a simple explicit solution is found as in (48), where $v_{e}=\sqrt{||u_{\text{unc}}||_{2}^{2}-||n^{\top}u_{\text{unc}}||_{2}^{2}}=||H^{\top}u_{\text{unc}}||_{2}$ is the projection of $u_{\text{unc}}$ onto the hyperplane $H=H(x,x_{o})$ tangent to the boundary function $h(x,t)$ (23).
@@ VI-C Theoretical Guarantees
Consider an obstacle whose boundary is defined as $h(x,x_{o})=0$. Any trajectory $x(t)$, $t\in[0,\infty)$ starting outside the obstacle boundary, i.e., $h(x(0),x_{o}(0)\geq 0$, evolving according to Eq. (21) and being constrained by the Speed-Constraining and Velocity-Constraining approaches Mod-DS presented in Section VI-A and VI-B, will never penetrate the obstacle boundary, i.e., $h(x(t),x_{o}(t))\geq 0$, $\forall t\in[0,\infty)$.
@@
VII Modulated Control Barrier Functions (MCBF)
In fully actuated systems, Mod-DS approaches outperform CBF-QP in navigating around non-convex unsafe sets with far fewer undesirable local minima. However, CBF-QP is still more commonly used in robot applications because it works for control affine systems in general and the QP formulations are more intuitive to the users. Nevertheless, choosing between superiority in handling complex robot dynamics and navigating among concave unsafe sets free of undesirable local minima should not be a tradeoff that users have to make when designing robot controllers. In this section, 2 Modulated CBF-QP (MCBF-QP) approaches that combine the strength of CBF-QP with Mod-DS are proposed to ensure respectively few-local-minimum and local-minimum-free navigation in concave-obstacle environments with the ease of incorporating different robot dynamics.
@@ VII-A Reference MCBF for Star-shaped Obstacle Avoidance
In subsection V-C, theoretical analysis of reference Mod-DS and CBF-QP demonstrated that CBF-QP’s performance in $n(x,x_{o})$ direction is similar to that of the reference Mod-DS in Theorem V.5. Besides, in Theorem V.6, reference Mod-DS’s introduction of off-diagonal components in the tangent hyperplane $H(x,x_{o})$ navigation is identified to be the major differences between CBF-QP and reference Mod-DS in fully-actuated systems and the key factor for reducing undesirable equilibrium in star-shaped obstacle avoidance. Adding additional constraints for $u_{\text{cbf}}$ on the tangent hyperplane will reduce undesirable equilibrium caused by CBF constraints. In theory, there are numerous ways to incorporate the off-diagonal components into CBF-QP results. Here, we propose a linear constraint formulation that is mathematically simple and does not increase the complexity of the original CBF-QP problem. Note that $E_{r}(x,x_{o})^{-1}[2:d,:]\dot{x}=H(x,x_{o})^{\top}(I-\frac{1}{\nabla_{x}h(x,x_{o})^{\top}r(x,x_{o})}r(x,x_{o})\nabla_{x}h(x,x_{o})^{\top})\dot{x}$ solves for the projection of $\dot{x}$ onto hyperplane $H(x,x_{o})$ in basis $E_{r}(x,x_{o})$. $g(x)$ and $E_{r}(x,x_{o})^{-1}$ are defined in (15) and (38).
@@
$L_{f}h(x,x_{o})+L_{g}h(x,x_{o})u+\nabla_{x_{o}}h(x,x_{o})\dot{x}_{o}\geq-\alpha(h(x,x_{o}))$ (50)
$H(x,x_{o})^{\top}(I-\frac{r(x,x_{o})\nabla_{x}h(x,x_{o})^{\top}}{\nabla_{x}h(x,x_{o})^{\top}r(x,x_{o})})g(x)(u-u_{\text{nom}})=\rho$ (51)
@@
$\nabla_{x}h(x,x_{o})^{\top}u+\nabla_{x_{o}}h(x,x_{o})\dot{x}_{o}\geq-\alpha(h(x,x_{o}))$ (53)
$H(x,x_{o})^{\top}(I-\frac{r(x,x_{o})\nabla_{x}h(x,x_{o})^{\top}}{\nabla_{x}h(x,x_{o})^{\top}r(x,x_{o})})(u-u_{\text{nom}})=\rho$ (54)
Reference-MCBF-QP, compared with regular CBF-QP, improves the performance of the optimization-based safe controller for star-shaped unsafe sets by reducing local minimum numbers. The new method ensures that the safe control outputs will never lead to undesirable equilibria, except when $\dot{x}_{\text{nom}}$ is inversely collinear with the reference direction $r(x,x_{o})$. The added constraints are designed to be of linear equality types and, thus would not cause large runtime increases. Additionally, the employment of relaxation parameter $\rho$ sustains the feasibility of the regular CBF-QP, i.e. given the same $x,x_{o}$, Reference-MCBF-QP will be feasible if and only if the CBF-QP without constraints (51) and (54) is feasible. The performance of fully-actuated reference MCBF-QP in star-shaped obstacle environments is validated in Figure 7.
@@
Derivation: To draw connections with Theorem V.6, here we again assume static-single-obstacle scenarios ($\dot{x}_{o}=\textbf{0}$) and define $\alpha=\alpha(h(x,x_{o}))$, $H=H(x,x_{o})$, $r=r(x,x_{o})$, and $\nabla_{x}h=\nabla_{x}h(x,x_{o})$. MCBF-QP in (52) can be reformulated into Lagrangian function in (55), where $\mu_{\text{eq}}\in\mathbb{R}^{d-1}$ and $\mu_{\text{cbf}}$ are the Lagrange multipliers.
@@
$V=\begin{cases}\begin{bmatrix}u_{\text{nom}}\\ 0\end{bmatrix}\quad\text{if}\quad\nabla_{x}h^{\top}u_{\text{nom}}+\alpha\geq 0\\ \begin{bmatrix}F^{-1}(I+\nabla_{x}hA)&-A^{\top}\\ A&-\frac{AF\nabla_{x}h}{\nabla_{x}h^{\top}\nabla_{x}h}\end{bmatrix}\begin{bmatrix}Fu_{\text{nom}}\\ -\alpha\end{bmatrix}\quad\text{otherwise}\end{cases},$ (59)
@@
In other words, the explicit solution $u_{\text{mcbf}}$ becomes
$u_{\text{mcbf}}=\begin{cases}u_{\text{nom}}\quad\text{if}\quad\nabla_{x}h^{\top}u_{\text{nom}}+\alpha\geq 0\\ F^{-1}(I+\nabla_{x}hA)Fu_{\text{nom}}+A^{\top}\alpha\quad\text{otherwise}\end{cases}.$ (61)
From 5(b), it is noticeable that vector $r(x,x_{o})$ can be represented as a weighted vector sum of $n(x,x_{o})$ and a vector $e_{i}(x,x_{o})$ respectively orthogonal and parallel to the hyperplane $H$, as in (62), $w_{i}\in\mathbb{R}^{+}$. Like before, we denote $e_{i}=e_{i}(x,x_{o})$ and $n=n(x,x_{o})$ to simplify the mathematical expressions. Note that $\sum_{i=0}^{d-1}(w_{i})^{2}=1$ because by definition, $||r||_{2}=1$.
@@
Based on the relationship between $n$ and $\nabla_{x}h$ given in (24), $\nabla_{x}h=||\nabla_{x}h||_{2}n$. Plug the new representation of $r$ and $\nabla_{x}h$ into (58), $F$ can be orthogonally diagonalized as the following, where $E=[n,H]$ is defined in Eq (23) and (24).
@@
Remark: The explicit solution of regular CBF-QP in fully-actuated systems (20) can be rewritten using $n$ as the following.
@@
Since $\lambda$, $\lambda_{e}$ can be any real number satisfying (25), there exist an infinite amount of $\lambda$, $\lambda_{e}$ meeting the requirement $\lambda_{e}-\lambda=\frac{1}{2}$. Therefore, the conclusion can be drawn that there always exist feasible $\lambda$, $\lambda_{e}$ pairs such that outputs from reference MCBF-QP equal that from reference Mod-DS on boundaries of the unsafe sets. This guarantees that reference MCBF-QP is free of undesirable equilibria except when $u_{\text{nom}}$ is not inversely collinear to $r(x,x_{o})$ in the boundary set $\partial C$.
@@ VII-B On-Manifold MCBF for Non-star-shaped Obstacle Mod- Avoidance
By keeping CBF constraints in regular CBF-QP to account for Mod-DS’s policies in $n(x,x_{o})$ direction and introducing new constraints (51) and (54) to mimic reference Mod-DS’s actions in the tangent hyperplane $H(x,x_{o})$, reference MCBF-QP successfully combines the strength of both and realizes star-shaped obstacle avoidance using QP-based methods. While on-manifold Modulation, functioning under the aggregate effects of 3 policies, is too structurally complicated to be explicitly connected to CBF-QP on the theory level, its main component, like other Mod-DS approaches, modulates $\dot{x}$ in $n(x,x_{o})$ direction similar to CBF constraints and prevents the formation of undesirable equilibrium by ordering the directions of $\dot{x}$ projected onto the tangent hyperplane $H(x,x_{o})$ using $\phi(x,x_{o})$. Therefore, by introducing $H(x,x_{o})$ related constraints, inspired by $\phi(x,x_{o})$ in on-manifold Mod-DS, into CBF-QP, on-manifold MCBF-QP is constructed to achieve local-minimum-free safe control for control affine systems in general (VII-B). The parameter $\gamma$ is a user-defined positive real number, i.e. $\gamma\in\mathbb{R}^{+}$. Unlike reference MCBF-QP, which stays feasible as long as the CBF-QP without constraint (51) is feasible, on-manifold MCBF-QP’s feasible domain is always smaller than that of the CBF-QP without constraint (75). Additionally, the larger $\gamma$ is, the smaller the feasible domain of on-manifold MCBF-QP will be. The performance of on-manifold MCBF-QP in concave obstacle environments is validated in Figure 8.
@@
$L_{f}h(x,x_{o})+L_{g}h(x,x_{o})u+\nabla_{x_{o}}h(x,x_{o})\dot{x}_{o}\geq-\alpha(h(x,x_{o}))$ (74)
$\phi(x,x_{o})^{\top}f(x)+\phi(x,x_{o})^{\top}g(x)u\geq\gamma$ (75)
@@
$\nabla_{x}h(x,x_{o})^{\top}u+\nabla_{x_{o}}h(x,x_{o})\dot{x}_{o}\geq-\alpha(h(x,x_{o}))$ (76)
$\phi(x,x_{o})^{\top}u\geq\gamma$ (77)
@@
Remark: While reference MCBF-QP can be in theory extended to nonlinear dynamical systems with equations in (VII-A), its applications in high dimensional systems is limited because finding reference directions $r(x,x_{o})$ when $d>3$ is nontrivial in general and its complexity scales up as $d$ increases. However, given proper step size $\beta$, the geodesic approximation method in (29) is able to closely approximate unsafe set boundaries for any given state $x\in\mathbb{R}^{d}$. As a result, on-manifold MCBF-QP can be readily applied in any control affine robot system given a reasonably designed barrier function $h(x,x_{o})$.